The volume of a cube is given by V= s to the 3rd power . Find the length of a side of a cube if the Volume is 729 cm to the 3rd power.
Answer:
Show work in this space.


This problem has me baffled as well.


Answer by funmath(2933)

The answer is: 9cm
The work:
s^3=729cm^3
3root(s^3)=3root(729cm^3)
3root(s^3)=3root(9*9*9cm^3)
3root(s^3)=3root(9^3cm^3)
s=9cm